OpenStudy (anonymous):

Step-fuctions: Evaluate the function for the indicated values: g(x)=2[[x]] a) g(-3) b) g(0.25) c) g(9.5) d) g(11/3) Wouldn't I plug in the greatest integer less than the value I'm plugging in? That's what my teacher said. So, for example, wouldn't I plug in -3 for a, 0 for b, 9 for c and 3 for d?

OpenStudy (anonymous):

@jim_thompson5910 can you help me please?

jimthompson5910 (jim_thompson5910):

[[x]] is the greatest integer function which means you round down to the nearest whole number examples: [[0.2]] = 0 [[6.8]] = 6 [[-2.1]] = -3 Basically wherever you are on the number line, you move left to the nearest integer. If you are already at an integer, you stay where you are.

jimthompson5910 (jim_thompson5910):

I'll do part b g(x)=2[[x]] g(0.25)=2[[0.25]] g(0.25)=2*0 g(0.25)=0

jimthompson5910 (jim_thompson5910):

so it looks like you have it correct

OpenStudy (anonymous):

Ok, so the answers would be: a) -3 b) 0 c) 9 d) 3 Correct? Because if that is, then I have one more question.

jimthompson5910 (jim_thompson5910):

you are forgetting about the 2 hanging out front

jimthompson5910 (jim_thompson5910):

the function is NOT g(x) = [[x]] it is g(x) = 2[[x]]

OpenStudy (anonymous):

Ohhh, so I have to multiply it by 2. Alright, got it. But I have one more question.

jimthompson5910 (jim_thompson5910):

ok go ahead

OpenStudy (anonymous):

The function is g(x)=3[[x-2]]+5, and the given values are: a) g(1/8) b) g(9) c) g(-4) d) g(3/2) I did the right thing but I'm still not getting the right answers. I know because I checked the answers in the back of my book but they're not correct.

jimthompson5910 (jim_thompson5910):

what answers are you getting

OpenStudy (anonymous):

Oops, I gave the wrong values for a-d. a) g(-2.7) b) g(-1) c) g(0.8) d) g(14.5) Sorry about that. And for the first one I got.... -7. The second one I got -4, the third one I got -1 and for the last one I got 41.

OpenStudy (anonymous):

Actually, all of them are right but the first one. Idk what I'm doing wrong in the first one.

jimthompson5910 (jim_thompson5910):

g(x)=3[[x-2]]+5 g(-2.7)=3[[-2.7-2]]+5 g(-2.7)=3[[-4.7]]+5 g(-2.7)=3*(-5)+5 ... round down g(-2.7)=-15+5 g(-2.7)=-10

jimthompson5910 (jim_thompson5910):

you do NOT do this g(x)=3[[x-2]]+5 g(-2.7)=3[[-2.7-2]]+5 g(-2.7)=3[[-4.7]]+5 g(-2.7)=3*(-4)+5 g(-2.7)=-12+5 g(-2.7)=-7 this is all wrong because -4.7 should round down to -5 and not to -4

OpenStudy (anonymous):

Ohhhh ok. So if I wanted to plug in the greatest integer less than -2.7, I'd plug in -3?

jimthompson5910 (jim_thompson5910):

that is if you had [[x]] and you plugged in x = -2.7 but you have [[x-2]]

jimthompson5910 (jim_thompson5910):

[[x]] means... wherever you are on the number line, you move left to the nearest integer. If you are already at an integer, you stay where you are.

OpenStudy (anonymous):

Right, so was I wrong about the -3? Cuz when I plugged in -3 for x in g(-2.7)=3[[-3-2]]+5, I got -10, which is correct.

jimthompson5910 (jim_thompson5910):

[[x]] is also known as the floor function because you go to the "floor", which basically the closest whole number

jimthompson5910 (jim_thompson5910):

well it is coincidental that [[-3-2]] = [[-2.7 - 2]] the error you made could produce a false answer down the road

jimthompson5910 (jim_thompson5910):

[[-3-2]] = [[-5]] = -5 [[-2.7-2]] = [[-4.7]] = -5 they lead to the same output, but they aren't the same

OpenStudy (anonymous):

Oh okay. That makes sense. I'll practice some more to get the hang of it. It's still a bit confusing. We just learned this today but I'll ask my teacher for help too.

jimthompson5910 (jim_thompson5910):

ok sounds good

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome